OpenStudy (anonymous):

Which equation represents the axis of symmetry of the function y = –2x² + 4x –6? (1 point) y = 1 x = 1 x = 3 x = –3

OpenStudy (anonymous):

y = –2x² + 4x –6 = x² - 2x +3 = (x-1)² + 4 so the vertex of the parabola is in (1,4) so answer is x= 1

OpenStudy (anonymous):

ups, made a mistake. is (x-1)² +2, but answr is the same anyway

OpenStudy (callisto):

even though your answer is correct for x=1, i think there's a problem with your steps. you cannot just cannel the '-2' y = –2x² + 4x –6 =-2( x² - 2x) -6 = -2(x-1)² - 4 vertex should be (1, -4) and the axis of symmetry is x=1

OpenStudy (anonymous):

i got 1 for x too